OpenStudy (anonymous):

Find the area of an equilateral triangle (regular 3-gon) with the given measurement. 6-inch radius A = sq. in. i have 18sqrt3but idk if its right or not please help me

OpenStudy (anonymous):

hmmm, well if it has a 6 inch radius, then that means that 6 inches is half of the height

OpenStudy (anonymous):

the area of an equilateral triangle is \[A=s ^{2}\sqrt{3}/4\]

OpenStudy (anonymous):

right, but what is the length of S?

OpenStudy (anonymous):

here is a good explanation about how to think about it: http://en.allexperts.com/q/Geometry-2060/area-equilateral-triangle-given.htm

OpenStudy (anonymous):

I'm guessing s=6 in this case since it's the only dimension we're given. As for your result, it seems close but using the given area for s=6, your answer would be \[9\sqrt{3}\]

OpenStudy (anonymous):

s=length of one side of the equalateral triangle, not the radius

OpenStudy (anonymous):

using trig though you can follow the link I pasted to get the length of s, given the radius dissects the angles in known ways

OpenStudy (anonymous):

no the radius would b half the height. there fore the base would b b=2(r/tan(30)) the height h=2r so A=1/2(2r)^2(1/tan(30))

OpenStudy (anonymous):

i already have 9sqrt3 for the side i needd the radius